A sum of money amounts to rs 5200 in 5 yeas and rs 5680 in 7 years at simple interest. What was the rate of simple interest charged per annum
Answer by jorel1380(3719)

Let P be the principal, and r be the interest rate. Then:
P(1+5r)=5200 and
P(1+7r)=5680
P+5Pr=5200
P+7Pr=5680
2Pr=480
Pr=240
r=240/P
P+1200=5200
P=4000
Solve for r. 🙂🙂🙂🙂